### v1

# MaNGA DynPop – VII. A Unified Bulge-Disk-Halo Model for Explaining Diversity in Circular Velocity Curves of 6000 Spiral and Early-Type Galaxies

Abstract
We derive circular velocity curves (CVCs) from stellar dynamical models for $\sim 6{,}000$ nearby galaxies in the final data release of the SDSS-IV MaNGA survey with integral-field spectroscopy, exploring connections between the inner gravitational potential (traced by CVC amplitude/shape) and galaxy properties. The maximum circular velocity ($V_{\rm circ}^{\rm max}$) and circular velocity at the half-light radius ($V_{\rm circ}(R_{\rm e}^{\rm maj})$) both scale linearly with the stellar second velocity moment $\sigma_{\rm e}^{2}\equiv\langle V^{2}+\sigma^{2}\rangle$ within the half-light isophote, following $V_{\rm circ}^{\rm max}\approx 1.75\sigma_{\rm e}$ (8% error) and $V_{\rm circ}(R_{\rm e}^{\rm maj})\approx 1.65\sigma_{\rm e}$ (9% error). CVC shapes (rising, flat, declining) correlate strongly with structural and stellar population properties: declining curves dominate in massive, early-type, bulge-dominated galaxies with old, metal-rich stars and early quenching, while rising CVCs prevail in disk-dominated systems with younger stellar populations and ongoing star formation. Using a unified bulge-disk-halo model, we predict CVC shapes with minimal bias, identifying three governing parameters: bulge-to-total mass ratio ($B/T$), dark matter fraction within $R_{\rm e}$, and bulge Sersic index. The distribution of CVC shapes across the mass-size plane reflects evolutionary pathways driven by (i) in situ star formation (spurring bulge growth) and (ii) dry mergers. This establishes CVC morphology as a diagnostic for galaxy evolution, linking dynamical signatures to structural and stellar population histories.
Galaxy dynamics(2667) — Galaxy formation(2669) — Galaxy evolution(2307) — Galaxy structure(2711) — Galaxy mass distribution(2703)
1 Introduction
The rotation curve (RC) of a galaxy represents the rotational velocity of stars or gas as a function of radial distance from the galaxy’s center. The first indirect evidence of dark matter (DM) came from studies of the velocity dispersion of galaxies within the Coma Cluster (Zwicky, 1933, 1937, 2009). Subsequent investigations into DM using galaxy RCs began with nearby galaxies. Babcock (1939) proposed that the nearly flat RC observed at the outskirts of M31 could be explained by an increasing mass-to-light ratio in the outer regions, while Oort (1940) identified substantial amounts of invisible matter in the outer parts of NGC 3115 based on its RC. With the improvements of observational instruments, the rotation curve (RC) has long been a powerful tool for studying the mass distributions of spiral galaxies, either using optical emission lines (Rubin & Ford, 1970; Rubin et al., 1980, 1982, 1985) or radio 21-cm line (Bosma & van der Kruit, 1979; Bosma, 1981).
As indicators of the gravitational potential (encompassing both baryonic and DM mass distributions) within which galaxies reside, the amplitude and the shape of RCs are expected to correlate with various galaxy properties such as luminosity, morphology, and structure. For instance, the Tully-Fisher relation (Tully & Fisher, 1977) tightly links the maximum rotation velocities (the amplitudes) and the total luminosities (or stellar masses) of spiral galaxies. Beyond this well-known correlation between RC amplitude and mass, the shapes of RCs have also been found to be dependent on galaxy morphologies (Corradi & Capaccioli, 1990; Erroz-Ferrer et al., 2016), luminosities (Persic et al., 1996; Sofue & Rubin, 2001), or a combination of both (Casertano & van Gorkom, 1991; Noordermeer et al., 2007; Swaters et al., 2009; Yoon et al., 2021). These evident correlations imply that the gravitational potential plays a crucial role in the evolution of galaxy structure and morphology.
In addition, the gravitational potential also has a significant impact on star-formation quenching, leaving imprints on the present-day stellar population properties. Various mechanisms that halt star formation of galaxies can be categorized into external (outside the galaxies) and internal (within the galaxies). The external mechanisms include ram pressure stripping (Gunn & Gott, 1972), strangulation (Larson et al., 1980), tidal stripping (Wetzel et al., 2013), dynamical heating from galaxy harassment (Moore et al., 1996), while internal mechanisms include but are not limited to halo quenching (Dekel & Birnboim, 2006), active galactic nucleus (AGN) feedback (Harrison, 2017), stellar feedback (Colling et al., 2018), mass quenching (Peng et al., 2010), bar quenching (Khoperskov et al., 2018), morphological quenching (Martig et al., 2009), and angular momentum quenching (Lu et al., 2022). Among these quenching mechanisms, the morphological quenching (Martig et al., 2009), also known as gravitational quenching (Genzel et al., 2014) or dynamical suppression (Davis et al., 2014; Gensior et al., 2020; Gensior & Kruijssen, 2021), occurs when the growth of a stellar spheroid or bulge stabilizes the gas disk instability (Toomre, 1964), thereby quenching star formation. This gravitational potential-related mechanism does not require gas removal (e.g. the tidal/ram pressure stripping) or massive halo (e.g. AGN feedback, shock heating induced halo quenching), and thus can explain the appearance of gas-rich but quenched galaxies in halos less massive than $10^{12}{\rm M_{\odot}}$. Linking the gravitational potential and stellar population properties to distinguish gravitational quenching from other quenching mechanisms is crucial in understanding the complex star-formation histories that shapes the evolution of galaxies.
The RCs were initially derived from long-slit spectroscopy and only provided one-dimensional (1D) information on stellar kinematics or ionized gas kinematics. Due to random or systematic misalignment of the slit (i.e. if the slit is not aligned with the galaxy’s major axis), there might be an offset in the measured RCs, even if the galaxy is perfectly axisymmetric. The advent of integral field unit (IFU) surveys, e.g. SAURON (de Zeeuw et al., 2002), ATLAS3D (Cappellari et al., 2011), CALIFA (Sánchez et al., 2012), SAMI (Bryant et al., 2015), and MaNGA (Bundy et al., 2015), provides two-dimensional (2D) kinematic information, reducing the effect of slit misalignment when extracting RCs (e.g. Yoon et al., 2021; Ristea et al., 2024a).
However, it was also pointed out that the RCs only work well in spirals and will underestimate the true CVCs of dispersion-dominated galaxies due to their significant non-circular motions (Roper et al., 2023; Downing & Oman, 2023; Sands et al., 2024). One way to overcome this intrinsic limitation of RCs and extend the studies to include early-type galaxies (ETGs) is applying corrections to account for the contribution of disordered motions, e.g. the asymmetric drift correction (Shetty et al., 2020; Bershady et al., 2024) or computing circular velocities from kinetic energy (Ristea et al., 2024b). A more direct and more accurate way is to constrain the gravitational potential (or CVCs) through detailed stellar dynamical models combined with spatially resolved stellar kinematics (e.g., Cappellari et al., 2013a; Li et al., 2017; Leung et al., 2018; Zhu et al., 2023b). CVCs derived from stellar dynamical models also have advantages over directly measured RCs, such as accounting for seeing effects and using accurately recovered inclination angle (Cappellari, 2008).
In this paper, which is the seventh paper of our DynPop series, we derive the CVCs from the JAM models for 10000 galaxies (of which 6000 are deemed reliable) provided in the MaNGA DynPop project (Paper I; Zhu et al., 2023b). The MaNGA DynPop project not only provides the mass distributions but also the stellar population properties (Paper II; Lu et al., 2023), enabling us to study the dynamical scaling relations in conjunction with stellar population properties (Paper III; Zhu et al., 2024), the density profiles for galaxy groups and clusters by combining stellar dynamics and weak lensing (Paper IV; Wang et al., 2024), the variation of stellar initial mass function (IMF) in ETGs (Paper V; Lu et al., 2024), and a detailed comparison of total density slopes between MaNGA and simulations (Paper VI; Li et al., 2024). The structure of this paper is organized as follows. Section 2 briefly introduces the MaNGA data and how we derive the CVCs and other galaxy properties. In Section 3, we study the correlations between CVCs (including the amplitude and the shape) and other galaxy properties (Section 3.1 and Section 3.2), propose a model to quantify the shape of CVCs (Section 3.3), and investigate the evolution of CVC shapes on the mass-size plane (Section 3.4). We summarize our main results in Section 4. Throughout the paper, we assume a flat Universe with $\Omega_{\rm m}=0.307$ and $H_{0}=67.7\,\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$ (Planck Collaboration et al., 2016), for consistency with the other papers of the DynPop series.
2 Sample and Data
2.1 The MaNGA survey
As one of the three projects in Sloan Digital Sky Survey-IV (SDSS-IV; Blanton et al., 2017), the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) survey (Bundy et al., 2015) provides spatially resolved spectral measurements of $\sim 10000$ nearby galaxies. The MaNGA project uses the IFU technique to obtain spectra simultaneously across the face of target galaxies, employing tightly-packed fiber bundles that feed into the BOSS spectrographs (Smee et al., 2013; Drory et al., 2015) on the Sloan 2.5m telescope (Gunn et al., 2006). The field-of-view (FoV) of MaNGA observations extends radially out to 1.5 effective radii ($R_{\rm e}$) for about two-thirds of the galaxies (Primary+ sample), and up to 2.5 $R_{\rm e}$ for roughly one-third of the galaxies (Secondary sample) at higher redshifts (Law et al., 2015; Wake et al., 2017). The final sample exhibits an approximately flat stellar mass distribution across the range $10^{9}-6\times 10^{11}\,\rm M_{\odot}$ (Wake et al., 2017), with a median redshift of $z\sim 0.03$.
The MaNGA spectra cover a wavelength range of $3600-10300\,\rm{\AA}$, with a spectral resolution of $\sigma=72\,\rm{\rm km\,s^{-1}}$ (Law et al., 2016). The raw data are spectrophotometrically calibrated (Yan et al., 2016) and processed using the Data Reduction Pipeline (DRP; Law et al., 2016) to produce data cubes. Stellar kinematic maps are then extracted from these data cubes through the Data Analysis Pipeline (DAP; Belfiore et al., 2019; Westfall et al., 2019), which utilizes the ppxf software (Cappellari & Emsellem, 2004; Cappellari, 2017, 2023) and a subset of the MILES stellar library (Sánchez-Blázquez et al., 2006; Falcón-Barroso et al., 2011), MILES-HC, to fit the absorption lines in the IFU spectra. Before extracting stellar kinematics, the spectra are Voronoi binned (Cappellari & Copin, 2003) to a signal-to-noise ratio of $\rm S/N=10$ to obtain reliable measurements.
2.2 Dynamical models, CVCs, and other dynamical properties
We performed Jeans Anisotropic Modelling (JAM; Cappellari, 2008, 2020) to construct dynamical models for the whole MaNGA sample in Zhu et al. (2023b). The JAM model allows for anisotropy in second velocity moments and two different assumptions on the orientation of velocity ellipsoid, i.e. $\rm JAM_{cyl}$ (cylindrically-aligned) and $\rm JAM_{sph}$ (spherically-aligned). Four different mass models are adopted in Zhu et al. (2023b), which differ primarily in their assumptions about dark matter distributions: (i) the mass-follows-light model which assumes that the total mass density traces the luminosity density (hereafter MFL model), (ii) the model which assumes a spherical NFW (Navarro et al., 1996) dark halo (hereafter NFW model), (iii) the fixed NFW model which assumes a spherical NFW halo predicted by the stellar mass-to-halo mass relation in Moster et al. (2013) and mass-concentration relation in Dutton & Macciò (2014) (hereafter fixed NFW model), (iv) the model which assumes a generalized NFW (Wyithe et al., 2001) dark halo (hereafter gNFW model). Further details about the eight models can be found in Zhu et al. (2023b), while the data catalog is available online. [^1] https://manga-dynpop.github.io/.
Based on comparisons of observed and modeled stellar kinematics, the entire sample is classified according to different modeling qualities (Qual = $-1$, 0, 1, 2, 3 from worst to best). In this work, we select 6065 galaxies that are flagged as $\rm Qual\geqslant 1$, for which the dynamical quantities related to the total mass distribution are nearly insensitive to variations in model assumptions (Zhu et al., 2023b). Throughout this paper, we adopt the gNFW model, which is the most flexible mass model, with the $\rm JAM_{cyl}$ assumption (the results and conclusions remain consistent when adopting the $\rm JAM_{sph}$ assumption) unless stated otherwise. The total mass distribution consists of three components: the nuclear supermassive black hole, the stellar mass distribution, and the dark matter mass distribution (i.e. a gNFW dark halo in this model). The black hole mass is estimated from $M_{\rm BH}-\sigma_{\rm c}$ relation (McConnell et al., 2011), where $\sigma_{\rm c}$ is computed as the mean stellar velocity dispersion within 1 FWHM of MaNGA PSF. For the stellar component, we use the Multi-Gaussian Expansion (MGE; Emsellem et al., 1994; Cappellari, 2002) method to fit SDSS r-band images and obtain the surface brightness. Then the surface brightness is deprojected to obtain the luminosity density of the kinematic tracer in the three-dimensional space. The gNFW profile is written as
$\rho_{{}_{\rm DM}}(r)=\rho_{s}\left(\frac{r}{r_{s}}\right)^{\gamma}\left(\frac{1}{2}+\frac{1}{2}\frac{r}{r_{s}}\right)^{-\gamma-3},$ (1)
where $r_{s}$ is the characteristic radius, $\rho_{s}$ is the characteristic density, and $\gamma$ is the inner density slope. For $\gamma=-1$, this function reduces to the NFW profile.
With the gravitational potential $\Phi(R,z)$ expressed in the form of MGE, we use the mge_vcirc procedure (equation 45 in Cappellari, 2020) in the jampy package to calculate the circular velocity $V_{\rm circ}$ at a given galactocentric radius $R$ in the equatorial plane. Assuming that the galaxies are axisymmetric, we further derive the maps with constant values of $V_{\rm circ}$ at a given galactocentric radius $R=\sqrt{x_{\rm bin}^{2}+y_{\rm bin}^{2}/q^{2}}$, where ($x_{\rm bin}$, $y_{\rm bin}$) are the coordinates of Voronoi bins and $q\equiv b/a$ is the axial ratio of half-light elliptical isophote derived from MGE. In Figure 1, we present examples of CVCs, circular velocity maps, line-of-sight velocity maps, stellar age maps and stellar metallicity maps for galaxies with different CVC shapes (rising, flat, or declining in the outskirts). Table 1 presents the parameters of CVCs, including the circular velocity at half-light radius and the maximum circular velocity within kinematic data range.
Table 1: Parameters of circular velocity curves for 6000 nearby galaxies.
plateifu mangaid DA $R_{\rm e}$ $R_{\rm e}^{\rm maj}$ $\rm rFWHM\_IFU$ $R(V_{\rm circ}^{\rm max})$ rmax $V_{\rm circ}(R_{\rm e})$ $V_{\rm circ}(R_{\rm e}^{\rm maj})$ $V_{\rm circ}^{\rm max}$ $V_{\rm circ}({\rm rmax})$ $\lg\,M_{\rm BH}$ Qual
[Mpc] [arcsec] [arcsec] [arcsec] [arcsec] [arcsec] [$\rm km~{}s^{-1}$] [$\rm km~{}s^{-1}$] [$\rm km~{}s^{-1}$] [$\rm km~{}s^{-1}$] [$\rm M_{\odot}$]
7443-1901 12-84620 81.139 4.291 4.789 2.580 5.750 5.750 108.584 116.479 130.848 130.848 4.580 0
7443-6103 12-84665 79.374 5.769 8.737 2.557 10.332 10.332 112.148 141.332 152.934 152.934 5.814 0
7443-3702 12-84670 428.710 3.295 3.298 2.540 2.674 6.456 404.585 404.550 411.914 377.347 8.389 2
7443-1902 12-49536 81.841 3.566 4.766 2.573 5.885 5.885 91.468 97.980 102.546 102.546 5.217 0
7443-9101 12-84660 170.514 6.951 7.461 2.591 11.591 11.591 109.817 110.667 121.424 121.424 5.911 0
7443-12702 12-84674 237.488 9.070 9.253 2.561 9.504 9.504 111.012 111.123 111.241 111.241 5.758 0
7443-12704 12-84731 81.049 13.045 26.411 2.537 3.334 14.369 131.025 159.826 141.175 133.302 6.388 3
7443-6102 12-180432 119.823 7.328 8.898 2.558 10.315 10.315 190.581 193.384 194.786 194.786 6.635 3
7443-3701 12-193534 77.483 4.717 6.054 2.596 6.285 7.374 80.656 82.244 82.247 81.593 5.309 0
7443-12701 12-98126 88.386 4.756 5.825 2.581 9.437 9.437 118.948 119.206 120.597 120.597 5.810 1
Note. — (1) plateifu: The name of plate ID + IFU design ID (e.g. 7443-1901); (2) mangaid: Unique MaNGA ID (e.g. 12-84620); (3) DA: Angular diameter distance which assumes a flat Universe with $\Omega_{\rm m}=0.307$ and $h=0.677$ (Planck Collaboration et al., 2016); (4) $R_{\rm e}$: Effective radius (projected circular half-light radius from MGE fitting, in SDSS r band); (5) $R_{\rm e}^{\rm maj}$: Major axis of elliptical half-light isophote from MGE fitting, in SDSS r band; (6) $\rm rFWHM\_IFU$: The PSF FWHM values of IFU observations, in SDSS r band; (7) $R(V_{\rm circ}^{\rm max})$: The galactocentric radius where has the maximum circular velocity; (8) rmax: The kinematic data range, which is defined as the largest radius of the Voronoi bins; (9) $V_{\rm circ}(R_{\rm e})$: Circular velocity at half-light radius in the equatorial plane; (10) $V_{\rm circ}^{\rm maj}$: Circular velocity at the major axis of half-light ellipse in the equatorial plane; (11) $V_{\rm circ}^{\rm max}$: The maximum circular velocity within kinematic data range; (12) $V_{\rm circ}({\rm rmax})$: The circular velocity at rmax. (13) $\lg\,M_{\rm BH}$: The mass of central super massive black hole adopted in our dynamical models, which is derived from scaling relation (see Section 2.2); (14) Qual: The quality of JAM models ($-1$ to 3 from worst to best), only galaxies with $\rm Qual\geqslant 1$ have reliable CVC measurements.
The full catalog is available as a machine-readable table.
Figure 1: Examples of different shapes of JAM-derived CVCs (rising CVCs for the first row, flat CVCs for the second row, and declining CVCs for the third row). For each galaxy, the RGB image with plateifu (top left), the map of circular velocities (top middle), the map of line-of-sight velocity (top right), the CVC in the equatorial plane derived from JAM model (bottom left), the map of stellar age (bottom middle), the map of stellar metallicity (bottom right) are presented from left to right and top to bottom. The black contours are the observed surface brightness contours in steps of 1 mag. The black dots are the centroids of the Voronoi bins from which the maps were linearly interpolated.
Other structural and dynamical quantities used in this work are mainly taken from Zhu et al. (2023b), we briefly introduce the parameters and the corresponding keywords here:
  • Ellipticity $\varepsilon$ (keyword: Eps_MGE): ellipticity of the half-light elliptical isophote derived from the MGE model.
  • Size parameters $R_{\rm e}$ (keyword: Re_arcsec_MGE) and $R_{\rm e}^{\rm maj}$ (keyword: Rmaj_arcsec_MGE): $R_{\rm e}$ is the circularized half-light radius (effective radius), while $R_{\rm e}^{\rm maj}$ is the semi-major axis of the half-light elliptical isophote. Both $R_{\rm e}$ and $R_{\rm e}^{\rm maj}$ in the catalog of Zhu et al. (2023b) have been scaled by a factor of 1.35 following Cappellari et al. (2013a).
  • Total luminosity in SDSS r-band $L$ (keyword: Lum_tot_MGE): $L$ is derived from the MGE model of the SDSS r-band image and has been corrected for the dust extinction (see more details about dust extinction in Lu et al., 2023).
  • Effective stellar velocity dispersion $\sigma_{\rm e}$ (keyword: Sigma_Re): the second moment of the line-of-sight velocity within the elliptical half-light isophote (with an area of $\pi R_{\rm e}^{2}$), defined as
    $\sigma_{\rm e}\equiv\langle V_{\rm rms}^{2}\rangle_{\rm e}^{1/2}\approx\sqrt{\frac{\sum_{k}F_{k}(V_{k}^{2}+\sigma_{k}^{2})}{\sum_{k}F_{k}}},$ (2)
    where $F_{k}$, $V_{k}$, and $\sigma_{k}$ are the flux, stellar velocity, and stellar velocity dispersion in the $k$-th IFU spaxel. This quantity closely approximates the velocity dispersion $\sigma$ that one would measure by co-adding all spectra within the same aperture, and fitting $(V,\sigma)$ for that spectrum (Cappellari et al., 2013a, section 4.3).
  • Stellar angular momentum proxy or spin parameter $\lambda_{\rm R_{e}}$ (keyword: Lambda_Re): $\lambda_{\rm R_{e}}$ is defined within the same aperture as $\sigma_{\rm e}$ (i.e. the elliptical half-light isophote), written as (Emsellem et al., 2007)
    $\lambda_{R_{\rm e}}=\frac{\sum_{k}F_{k}R_{k}|V_{k}|}{\sum_{k}F_{k}R_{k}\sqrt{V_{k}^{2}+\sigma_{k}^{2}}},$ (3)
    where $F_{k}$, $V_{k}$ and $\sigma_{k}$ are the same as Equation (missing) 2; $R_{k}$ is the distance of $k$-th spaxel to the galaxy centre. The $\lambda_{\rm R_{e}}$ in the catalog of Zhu et al. (2023b) has been corrected for the beam smearing effect following Graham et al. (2018) [^2] https://github.com/marktgraham/lambdaR_e_calc.
  • Dynamical mass $M_{\rm JAM}$: $M_{\rm JAM}$ is defined as
    $M_{\rm JAM}\equiv(M/L)_{\rm JAM}\times L,$ (4)
    where $(M/L)_{\rm JAM}$ is the dynamical mass-to-light ratio (keyword: log_ML_dyn) from the $\rm JAM_{cyl}$+MFL model and $L$ is the SDSS r-band total luminosity (keyword: Lum_tot_MGE).
[…]
  • The ratio between scale radius of NFW halo and luminous half-light radius $r_{\rm s}/R_{\rm e}$: $r_{\rm s}$ is estimated from the Chabrier IMF-based stellar mass, which is converted from the Salpeter IMF-based stellar mass $M_{\rm\ast}$ (see the definition in Section 2.3) by subtracting 0.215 dex (Madau & Dickinson, 2014, figure 4), the stellar mass-to-halo mass relation in Moster et al. (2013), and the mass-concentration relation in Dutton & Macciò (2014).
2.3 Stellar population properties based on Stellar Population Synthesis (SPS)
The stellar population properties used in this work come from Lu et al. (2023). They fit the IFU spectra of the MaNGA DRP (Law et al., 2016) data cubes using the ppxf software (Cappellari & Emsellem, 2004; Cappellari, 2017, 2023) with the fsps models (Conroy et al., 2009; Conroy & Gunn, 2010), the Salpeter (Salpeter, 1955) IMF, and the MIST isochrones (Choi et al., 2016). The properties are briefly summarized below:
  • Luminosity-weighted stellar age $\langle\lg\rm Age\rangle$ (keyword: LW_Age_Re), luminosity-weighted stellar metallicity $\langle[Z/H]\rangle$ (keyword: LW_Metal_Re) and their radial gradients (keywords: LW_Age_Slope and LW_Metal_Slope): the luminosity-weighted values are calculated as
    $\langle x\rangle\equiv\frac{\sum^{N}_{i=1}w_{i}L_{i}x_{i}}{\sum^{N}_{i=1}w_{i}L_{i}},$ (5)
    where $w_{i}$ is the fraction of mass contributed by the $i$-th template (the ppxf fit weight, if the templates are normalized to unitary mass), $L_{i}$ is the SDSS r-band luminosity per unit mass of the $i$-th template, and $x_{i}$ is the $\rm\lg Age$ (or $\mathrm{[Z/H]}$) of the $i$-th template. The radial gradients are measured by linearly fitting the $\langle\lg\rm Age\rangle$ or $\langle[Z/H]\rangle$ profiles within the elliptical half-light isophote.
  • Star formation histories $T_{50}$ (keyword: T50) and $T_{90}$ (keyword: T90): $T_{50}$ and $T_{90}$ are defined as the cosmic times when the stars that account for 50% and 90% of galaxies’ present-day stellar mass are formed, respectively. Note that $T_{50}$ and $T_{90}$ are given as lookback times in the catalog of Lu et al. (2023) and here we convert them into cosmic time based on the cosmological parameters used in this work.
  • Averaged intrinsic stellar mass-to-light ratio $(M_{\ast}/L)_{\rm SPS}$ (keyword: ML_int_Re) within the elliptical half-light isophote, which is calculated as
    $(M_{\ast}/L)_{\rm SPS}=\frac{\sum_{i=1}^{N}w_{i}M_{i}^{\rm nogas}}{\sum_{i=1}^{N}w_{i}L_{i}},$ (6)
    where $M_{i}^{\rm nogas}$ is the stellar mass of the $i$-th template, which includes the mass of living stars and stellar remnants but excludes the mass of lost gas during stellar evolution.
  • Total stellar mass $M_{\ast}$ defined as
    $M_{\ast}=(M_{\ast}/L)_{\rm SPS}\times L,$ (7)
    where $(M_{\ast}/L)_{\rm SPS}$ is the SDSS r-band stellar mass-to-light ratio derived from the stacked spectrum within the elliptical half-light isophote (keyword: ML_int_Re), $L$ is the total luminosity derived from MGE model (keyword: Lum_tot_MGE).
3 The circular velocity curves across various galaxy types
[…]
3.1 The amplitude of CVCs
The amplitude of CVCs is usually characterized by the maximum (asymptotic) circular velocity (Tully & Fisher, 1977) for spirals, while the characteristic velocity measurement for ETGs is typically the velocity dispersion which enters the Faber-Jackson relation (Faber & Jackson, 1976) and Fundamental Plane (Djorgovski & Davis, 1987; Dressler et al., 1987). Previous studies aimed at unifying the dynamical relations of spirals and ETGs were done by converting the velocity dispersion $\sigma_{\rm e}$ into the circular velocity at half-light radius $V_{\rm circ}(R_{\rm e}^{\rm maj})$ or the maximum circular velocity $V_{\rm circ}^{\rm max}$ (Padmanabhan et al., 2004; Courteau et al., 2007; Schulz et al., 2010; Dutton et al., 2011; Cappellari et al., 2013a). Here, we revisit the empirical relations (in the form $V_{\rm circ}=k\sigma_{\rm e}$, where $k$ is the conversion factor) using the accurate $\sigma_{\rm e}$ provided by MaNGA (Law et al., 2021), together with the circular velocities derived from our dynamical models.
The correlation between $\sigma_{\rm e}$ and $V_{\rm circ}(R_{\rm e}^{\rm maj})$ in MaNGA is presented in the top left panel of Figure 2. The best-fitting relation is given by
$\lg V_{\rm circ}(R_{\rm e}^{\rm maj})=2.36+0.87\times(\lg\sigma_{\rm e}-2.11),$ (8)
or approximately equals to
$V_{\rm circ}(R_{\rm e}^{\rm maj})\approx 3.31\times\sigma_{\rm e}^{0.87}\approx 1.65\times\sigma_{\rm e},$ (9)
with an error of $\Delta/\sqrt{2}=9$%. When classifying the sample into different morphological types following Zhu et al. (2024, section 2.5), the conversion factors for ETGs ($k\approx 1.63$) and LTGs ($k\approx 1.71$) are similar. Comparisons with previous observation (ATLAS3D; Cappellari et al., 2013a), the EAGLE (Schaye et al., 2015; Crain et al., 2015) and Illustris-TNG (Naiman et al., 2018; Springel et al., 2018; Pillepich et al., 2018; Marinacci et al., 2018; Nelson et al., 2018) cosmological simulations (the relations taken from Ferrero et al., 2021), and theoretical prediction (Wolf et al., 2010) are shown in the top right panel. Cappellari et al. (2013a) found a slightly smaller factor $k\approx 1.51$ in ETGs, while Wolf et al. (2010) predicted $k=\sqrt{3}$. The relation in the EAGLE simulations also has a similar factor of $k\approx 1.72$ (or equivalently $V_{\rm circ}(R_{\rm e})=2.37\times\sigma_{\rm e}^{0.94}$). However, unlike the relations mentioned above, which show weak $\sigma_{\rm e}$ dependence, the relation in Illustris-TNG simulations ($V_{\rm circ}(R_{\rm e})=7.03\times\sigma_{\rm e}^{0.73}$) has an exponent significantly deviating away from one.
A slightly tighter correlation between $\sigma_{\rm e}$ and $V_{\rm circ}^{\rm max}$ is observed in the bottom left panel of Figure 2, with
$\lg V_{\rm circ}^{\rm max}=2.37+0.93\times(\lg\sigma_{\rm e}-2.11),$ (10)
or approximately
$V_{\rm circ}^{\rm max}\approx 2.56\times\sigma_{\rm e}^{0.93}\approx 1.75\times\sigma_{\rm e},$ (11)
with an error of $\Delta/\sqrt{2}=8$%. No difference in the conversion factors are found for ETGs and LTGs (see the bottom left panel of Figure 2). This result is consistent with the conversion factor found for ETGs in ATLAS3D ($k\approx 1.76$; Cappellari et al., 2013a). Note that $V_{\rm circ}^{\rm max}$ is defined as the maximum circular velocity within the region where we have stellar kinematic data (usually 1.5 $R_{\rm e}$ or 2.5 $R_{\rm e}$ in MaNGA). When we only consider the galaxies whose radius of maximum circular velocity lies in the kinematic data range, the conversion factor $k\approx 1.75$ remains nearly unchanged but with a smaller error of $\Delta/\sqrt{2}=6$% (bottom right panel of Figure 2).
[…]
3.2 The shape of CVCs
In Figure 3, we present the normalized CVCs colored according to different galaxy properties. The top panels show a clear dependence of CVC shape on mass and morphology. Galaxies with higher mass (either dynamical mass $M_{\rm JAM}$ or stellar mass $M_{\ast}$), smaller T-Type values (indicating earlier types), and higher $B/T$ tend to have declining CVCs. In contrast, less massive, later-type galaxies with smaller $B/T$ tend to have rising CVCs. Flat CVCs at the outskirts are found in galaxies with intermediate mass and morphology (early-type spirals or S0), consistent with previous studies (Kalinova et al., 2017; Yoon et al., 2021).
The shape of CVCs also correlates with kinematic properties such as stellar velocity dispersion $\sigma_{\rm e}$ and the proxy for stellar angular momentum $\lambda_{\rm R_{e}}$ (first two panels in the second row). As $\sigma_{\rm e}$ increases and $\lambda_{\rm R_{e}}$ decreases, trends transition from rising CVCs to flat, and finally to declining CVCs. These trends reflect differences in galaxy structure: $\sigma_{\rm e}$ traces the bulge fraction (Cappellari et al., 2013b), while $\lambda_{\rm R_{e}}$ values reflect the fraction of ordered motions, which is also related to the prominence of the disk component or bulge mass fraction.
The $T_{50}$ and $T_{90}$ values represent the formation times of stars (in cosmic time) that account for 50% and 90% of the galaxy’s present-day stellar mass, respectively. The shape of CVCs shows clear dependence on $T_{90}$: declining CVCs for earlier quenched galaxies (higher $T_{90}$) and rising CVCs for later quenched galaxies (lower $T_{90}$). According to the two-phase evolutionary scenario on the mass-size plane (as discussed in Zhu et al., 2024), galaxies increase their mass and size through gas accretion-induced in situ star formation and dry mergers (by accreting ex situ already-formed stars). Earlier-quenched galaxies have already formed large bulges through in situ star formation and continue morphological transformation through dry mergers, while later-quenched galaxies still undergo bulge growth and transformation into earlier types. In contrast, there is little or no dependence on $T_{50}$ because $T_{50}$ is less relevant for distinguishing between the two mechanisms.
We also examine the dependence of CVC shape on stellar age and metallicity, as well as their gradients, in the bottom panels of Figure 3. Older and metal-rich galaxies with flatter age gradients tend to have declining CVCs. Conversely, rising CVCs usually appear in younger and metal-poor galaxies with negative age gradients. The weak dependence on metallicity gradients may relate to different trends of metallicity and metallicity gradients on the mass-size plane (see figures 8 and 12 in Lu et al., 2023): metallicity follows velocity dispersion well, but the distribution of metallicity gradients is more complex.
In summary, our main findings are as follows:
  1. (i)
    Galaxies with declining CVCs tend to be massive, early-type, early-quenched, old, metal-rich, with high $B/T$, high velocity dispersion, low spin, and flat age gradients.
  2. (ii)
    In contrast, galaxies with rising CVCs display opposite characteristics, while those with flat CVCs have features between declining and rising CVCs.
  3. (iii)
    There is a weak dependence on $T_{50}$ and metallicity gradients.
These quantities often relate to bulge and disk fractions, which are seen as the physical origins of the diverse CVC shapes. One might assume it is possible to predict CVC shape given bulge and disk fractions, and dark matter fraction that accounts for the dark matter halo contribution. In the next section, we will assess whether this model (with a bulge, a disk, and a dark matter halo) can reproduce the observed diverse CVC shapes.
Figure 3: The normalized CVCs correlate with other galaxy properties. Definitions can be found in Section 2.2 and Section 2.3. These properties include masses (dynamical mass $M_{\rm JAM}$, stellar mass $M_{\ast}$ based on Salpeter IMF), morphology (T-Type and $B/T$), velocity dispersion $\sigma_{\rm e}$, spin parameter $\lambda_{\rm R_{e}}$, the cosmic times when stars accounting for 50% ($T_{50}$) and 90% ($T_{90}$) of galaxies’ present-day stellar mass were formed, luminosity-weighted stellar age, stellar metallicity, and their radial gradients.
3.3 Bulge-disk-halo model to quantify the shape of CVCs
Based on the remarkable similarity of RCs first noted by Rubin et al. (1985), there have been some studies trying to predict the amplitude and shape of RCs with a few key photometric parameters. For instance, Persic et al. (1996) and Salucci et al. (2007) proposed a universal rotation curve (URC) model for spiral galaxies, and Karukes & Salucci (2017) and Di Paolo et al. (2019) extended the URC model to dwarf and low surface brightness galaxies, respectively. Recently, Patel et al. (2024) developed a neural network (NN) model to predict the shape of RCs using the same photometric quantities as the literature URC, and found higher accuracy with their NN-based URC. They concluded that the improvement of literature URC at all radii requires detailed modeling in the inner region (including a bulge component) and at the outskirts (better parameterization of dark matter halo). The inclusion of a bulge component is even more important when quantifying the shape of CVCs in MaNGA galaxies, which include a large fraction of bulge-dominated systems.
In this section, we try to use a bulge-disk-halo (hereafter BDH) model to predict the shape of CVCs. In this model, the bulge and disk components are described as two Sersic components, with the Sersic indices for the bulge and disk being $n_{\rm Ser,bulge}$ and $n_{\rm Ser,disk}=1$ (exponential), respectively, while the dark matter halo is assumed to be an NFW profile. For each Sersic component, the surface brightness is written as
$\Sigma(R)=\Sigma_{0}{\rm exp}\left[-b_{n}\left(\frac{R}{R_{e}}\right)^{1/n}\right],$ (12)
where $b_{n}$ is a function of Sersic index $n$, obtained by solving the equation (rewritten from Ciotti, 1991, equation 5)
$\frac{\Gamma(2n,b_{n})}{\Gamma(2n)}=Q(2n,b_{n})=\frac{1}{2},$ (13)
where $\Gamma(a)$ is the gamma function, $\Gamma(a,x)$ is the upper incomplete one, and $Q(a,x)$ is the regularized upper incomplete gamma function (Olver et al. 2010, equation 8.2.4). Various approximations have been proposed to calculate $b_{n}$ (e.g., Capaccioli, 1989; Prugniel & Simien, 1997; Ciotti & Bertin, 1999; Cappellari, 2023). However, one can also compute $b_{n}$ using the special function $Q^{-1}(a,s)$ giving the inverse of the regularized upper incomplete gamma function. This provides the solution for $z$ of $s=Q(a,z)$. From Equation (missing) 13 one can then just write
$b_{n}=Q^{-1}(2n,1/2),$ (14)
The function $Q^{-1}(a,s)$ is implemented in popular languages like Python as scipy.special.gammainccinv, Mathematica as InverseGammaRegularized, or Matlab as gammainccinv.
We adopt the analytical approximation in Lima Neto et al. (1999) to calculate the deprojected mass density of the Sersic model
$\rho(r)=\rho_{0}\left(\frac{r}{R_{e}}\right)^{-p_{n}}{\rm exp}\left[-b_{n}\left(\frac{r}{R_{e}}\right)^{1/n}\right],$ (15)
where $p_{n}$ is
$p_{n}=1-\frac{0.594}{n}+\frac{0.055}{n^{2}},$ (16)
and the enclosed mass profile
$M(r)=M_{\infty}\frac{\gamma[(3-p_{n})n,b_{n}(r/R_{e})^{1/n}]}{\Gamma[(3-p_{n})n]},$ (17)
where the enclosed mass at infinity (i.e. the total mass) is
$M_{\infty}=4\pi\rho_{0}R_{e}^{3}\,\frac{n\Gamma[(3-p_{n})n]}{b_{n}^{(3-p_{n})n}}.$ (18)
The readers are referred to Vitral & Mamon (2020) for a summary of different approximations for the deprojection of Sersic profile. For convenience, we assume a total stellar mass $M_{\rm\ast}$ and a projected half-light radius $R_{\rm e}$, which will finally be removed in the normalized CVC. We use the bulge-to-total ratio $B/T$, $R_{\rm e}$, and the $R_{\rm e,bulge}/R_{\rm e,disk}$ ratio to calculate $R_{\rm e,bulge}$ and $R_{\rm e,disk}$. Then we can obtain $M_{\rm\infty,bulge}=M_{\rm\ast}\times{\rm B/T}$ and $M_{\rm\infty,disk}=M_{\rm\ast}\times{\rm(1-B/T)}$ assuming spatially constant stellar mass-to-light ratio, and further the $M_{\rm bulge}(r)$ and $M_{\rm disk}(r)$ profiles from Equation (missing) 17. Unlike previous studies that adopt different disk shapes depending on morphological types (e.g. a spherical disk for ETGs and an infinitely thin disk for LTGs, as discussed in section 2.1 of Dutton et al., 2011), we assume a spherical disk for all galaxies. This assumption avoids introducing discontinuities by making the disk shape changing from spherical to infinitely thin depend on morphologies. Although the spherical disk is a crude approximation, our toy model remains meaningful as it adopts various fitting functions of model parameters and does not pretend to be quantitatively accurate. Moreover, adopting an alternative extreme assumption that all galaxies have infinitely thin disk (Freeman, 1970) leads to a similar level of accuracy in our model (see Appendix A and Figure 7), which justifies that the choice of a universal disk shape assumption does not significantly affect the results. For the NFW dark matter halo with two free parameters ($r_{\rm s}$ and $\rho_{\rm s}$), we use the $r_{\rm s}/R_{\rm e}$ value to estimate $r_{\rm s}$ and use the dark matter fraction within a sphere of $R_{\rm e}$, defined as
$f_{\rm DM}(<R_{\rm e})=\frac{M_{\rm DM}(R_{\rm e})}{M_{\rm bulge}(R_{\rm e})+M_{\rm disk}(R_{\rm e})+M_{\rm DM}(R_{\rm e})},$ (19)
to calculate the $\rho_{\rm s}$ value of the NFW profile and finally obtain $M_{\rm DM}(r)$. The circular velocity curve
$V_{\rm circ}(r)=\sqrt{\frac{G\times[M_{\rm bulge}(r)+M_{\rm disk}(r)+M_{\rm DM}(r)]}{r}}$ (20)
is normalized by $R_{\rm e}$ and $V_{\rm circ}(R_{\rm e})$ to remove the dependence on the assumed $M_{\ast}$ and $R_{\rm e}$. To summarize, the normalized CVCs, $V_{\rm circ}(R)/V_{\rm circ}(R_{\rm e})$ versus $R/R_{\rm e}$ [^3] Throughout this paper, we use the circular velocity in the equatorial plane ($z=0$), which means $r=\sqrt{R^{2}+z^{2}}=R$., can be predicted using the five free parameters: $B/T$, $R_{\rm e,bulge}/R_{\rm e,disk}$, $n_{\rm Ser,bulge}$, $r_{\rm s}/R_{\rm e}$, and $f_{\rm DM}(<R_{\rm e})$.
Rather than predicting the CVC shape for individual galaxies, our aim is to use this simple toy model to predict the overall trends for specific galaxy populations. We use the scaling relations shown in Figure 4, along with the $f_{\rm DM}(<R_{\rm e})-\sigma_{\rm e}$ relation in Lu et al. (2024, equation 7), to predict $B/T$, $n_{\rm Ser,bulge}$, $R_{\rm e,bulge}/R_{\rm e,disk}$, $r_{\rm s}/R_{\rm e}$, and $f_{\rm DM}(<R_{\rm e})$ for a given $\sigma_{\rm e}$. The best-fitting relations (red solid curves in Figure 4) for the median trends, assuming constant errors, are derived using arctan-based sigmoid functions:
$B/T=0.54-0.17\,\arctan\left[-16.37\,(\lg\sigma_{\rm e}-2.25)\right]$ (21)
$\lg n_{\rm Ser,bulge}=0.43+0.37\,\arctan\left[3.08\,(\lg\sigma_{\rm e}-2.19)\right]$ (22)
The $B/T-\lg\,\sigma_{\rm e}$, $\lg\,n_{\rm Ser,bulge}-\lg\,\sigma_{\rm e}$, and $f_{\rm DM}(<R_{\rm e})-\lg\,\sigma_{\rm e}$ (Lu et al., 2024, figure 2) relations show strong dependence on $\sigma_{\rm e}$, while $R_{\rm e,bulge}/R_{\rm e,disk}$ varies little with $\sigma_{\rm e}$. Although the median relation of $\lg\,(r_{\rm s}/R_{\rm e})-\lg\,\sigma_{\rm e}$ increases with higher $\sigma_{\rm e}$, however, the minimum $r_{\rm s}$ (for the median relation) is at least five times larger than $R_{\rm e}$, making the dark matter density profile within the kinematic data range is approximately power-law. Thus we use the median values $\lg\,(R_{\rm e,bulge}/R_{\rm e,disk})=-0.37$ and $\lg\,(r_{\rm s}/R_{\rm e})=0.74$ (the median value for $\lg\,(\sigma_{\rm e}/{\rm km\,s^{-1}})<2.1$, below which dark matter fractions become more significant) instead of the scaling relations for these two parameters.
Figure 5 shows the comparisons between CVC shapes predicted by the BDH model and those derived from JAM. As $\sigma_{\rm e}$ increases, galaxies’ $f_{\rm DM}(<R_{\rm e})$ decreases, $B/T$ increases, and $n_{\rm Ser,bulge}$ becomes higher, the shape of CVCs changes from rising (blue curve) to flat (cyan curve) and then declining (red curve). A remarkably small systematic bias can be seen for different $\sigma_{\rm e}$ bins (with a bin width of 0.1 dex), in which we compare the median profile (dashed) of JAM-derived CVCs and the BDH-predicted CVC (solid). This indicates that the physical origin of the diverse CVC shapes is the relative contribution (both the mass ratios and scale radii) of different components (reflected by the scaling relations above), and the dependence on stellar mass found in previous studies is merely due to those relations being dependent on stellar mass. Among these relations, the two most important are $B/T$ and $f_{\rm DM}(<R_{\rm e})$, while the $n_{\rm Ser,bulge}$ is less important (comparing Figure 5 and Figure 8).
An interesting fact is that, given a velocity dispersion $\sigma_{\rm e}$, we can estimate the amplitude $V_{\rm circ}(R_{\rm e}^{\rm maj})$ (or $V_{\rm circ}^{\rm max}$) of CVC with an error of 9% (or 8%) and predict the CVC shape with nearly zero bias in most cases. This demonstrates that $\sigma_{\rm e}$ is a good proxy for the inner gravitational potential. Especially when IFU observations for high-redshift galaxies are quite expensive, one can apply the empirical aperture correction for velocity dispersion (Zhu et al., 2023a) to obtain $\sigma_{\rm e}$ as the proxy for inner gravitational potential.
Figure 4: Scaling relations between structural parameters ($B/T$, $n_{\rm Ser,bulge}$, $R_{\rm e,bulge}/R_{\rm e,disk}$, $r_{\rm s}/R_{\rm e}$) and $\sigma_{\rm e}$ for the $\rm Qual\geqslant 1$ galaxies. $B/T$, $n_{\rm Ser,bulge}$, and $R_{\rm e,bulge}/R_{\rm e,disk}$ are taken from Domínguez Sánchez et al. (2022) (see Section 2.2 for definitions). $r_{\rm s}/R_{\rm e}$ is the ratio between scale radius of dark halo and $R_{\rm e}$ of the whole galaxy, where $r_{\rm s}$ is estimated from the Chabrier IMF-based stellar mass, the stellar mass-to-halo mass relation in Moster et al. (2013), and the mass-concentration relation in Dutton & Macciò (2014). The blue contours are the kernel density estimate for the galaxy distribution, while the black dashed curve and the gray shaded region represent the median value and [16th, 84th] percentile of values, respectively. The best-fitting relations to the median trends (Equation (missing) 21 and Equation (missing) 22) are shown in red solid curves. The horizontal red dashed line in the third panel represents the median value of $\lg\,(R_{\rm e,bulge}/R_{\rm e,disk})$ for the entire sample, while the horizontal red dashed line in the fourth panel represents the median value of $\lg\,(r_{\rm s}/R_{\rm e})$ for galaxies with $\lg\,(\sigma_{\rm e}/{\rm km\,s^{-1}})<2.1$.
Figure 5: Comparisons between the BDH model-predicted shape of normalized CVCs (solid curves) and those derived from JAM (dashed curves) for four different velocity dispersion ($\sigma_{\rm e}$) bins, indicated by different color. In the left panel, we show the median profile of the JAM-derived CVCs for each $\sigma_{\rm e}$ bin, while the BDH model predictions are calculated at the bin centers using the scaling relations (three relations as a function of $\sigma_{\rm e}$, and two are constants) presented in the left panel. The ratio between the JAM-derived and BDH-predicted CVCs, along with the associated error (represented by the [16th, 84th] percentiles of the JAM-derived CVCs), is shown in the right panel.
Figure 6: The distributions of $B/T$ (top left panel), $f_{\rm DM}(<R_{\rm e})$ (top right panel, taken from figure 16 in Zhu et al., 2024), $\lg\,n_{\rm Ser,bulge}$ (bottom left panel), and the median CVC within a given mass and size bin (bottom right panel) are shown on the $M_{\rm JAM}-R_{\rm e}^{\rm maj}$ plane. The distributions are smoothed by the loess software (Cappellari et al., 2013b) with frac=0.05. In all panels, the dashed lines correspond to effective velocity dispersion $\sigma_{\rm e}=$ 50, 100, 200, 300, 400, and 500 $\rm km\,s^{-1}$ from left to right, which are calculated using the scalar virial equation $\sigma_{\mathrm{e}}^{2}\equiv GM_{\rm JAM}/(5R_{\rm e}^{\rm maj})$ (with factor 5 from Cappellari et al. 2006). The magenta curve shows the zone of exclusion (ZOE) defined in (Cappellari et al., 2013b, equation 4), with the ZOE above $M_{\rm JAM}=2\times 10^{10}{\rm M_{\odot}}$ is approximately $R_{\rm e}^{\rm maj}\propto M_{\rm JAM}^{0.75}$. The gray contours show the kernel density estimate for the galaxy distribution. Bottom right panel: The inset panels in the bottom right panel show the median (red solid) and the [16th, 84th] percentiles (red shaded region), as well as the median FWHM of PSF (vertical shaded gray region). All the inset panels have the same physical scales as the bottom right panel, while the ticks and labels are shown in the upper left inset panel.
3.4 CVC shape on the mass-size plane
We revisit galaxy evolution (focusing on the CVC shapes) on the mass-size plane in Figure 6, in which we show the distributions of $B/T$, $f_{\rm DM}(<R_{\rm e})$, $n_{\rm Ser,bulge}$, and the median CVC for galaxies in each mass-size bin (corresponding to each inset panel). The bottom right panel reveals clear trends:
  • Moving from left to right on the mass-size plane, the CVC shapes transition from rising in lower-mass galaxies to declining in higher-mass galaxies.
  • At fixed mass, galaxies with larger size tend to have rising CVC shapes, while smaller galaxies typically exhibit declining CVCs.
  • Galaxies closest and parallel to the zone of exclusion (ZOE; $R_{\rm e}^{\rm maj}\propto M_{\rm JAM}^{0.75}$), which have the highest velocity dispersion of $\sigma_{\rm e}\approx 200\,{\rm km\,s^{-1}}$ (see the rightmost three inset panels), maintain declining CVCs, with little change in shape as mass and size increase.
As discussed in previous sections, the shape of CVCs is determined by $B/T$, $f_{\rm DM}(<R_{\rm e})$, and $n_{\rm Ser,bulge}$. Consequently, the evolution of CVC shapes can be understood through the evolution of these parameters. As suggested by the IFU results (Cappellari et al., 2013b; Cappellari, 2016; Zhu et al., 2024) and the observations of high-redshift ETGs (van der Wel et al., 2008; van Dokkum et al., 2015; Derkenne et al., 2021), galaxy evolution on the mass-size plane can be explained by the two-phase evolutionary scenario: (i) gas accretion-induced star formation and (ii) dry mergers.
When disk galaxies form (i.e., the formation of stellar disks; Mo et al., 1998), they typically exhibit small bulges (low $\sigma_{\rm e}$ and low $B/T$). These young disk galaxies tend to host pseudobulges (e.g. Hu et al., 2024), which typically exhibit exponential-like profiles with Sersic indices of $n_{\rm Ser,bulge}\approx 1-2$ (Kormendy & Kennicutt, 2004). Subsequent inner star formation triggered by gas accretion, drives bulge growth, resulting in a slightly higher Sersic index for the bulge and a lower central dark matter fraction. This transformation causes galaxies to shift their CVC shapes from rising to declining as their bulges and $\sigma_{\rm e}$ increase. The size dependence (quantified by the half-light radius) is primarily explained by the fact that, at fixed mass, galaxies with higher $B/T$ are smaller, and contain less dark matter within their smaller size (lower dark matter fraction). The trend differs for high $\sigma_{\rm e}$ ETGs that lie parallel to the ZOE, where dry mergers dominate mass and size growth. Along the direction of the ZOE, dry mergers mildly increase $B/T$ and have minimal impact on $f_{\rm DM}(<R_{\rm e})$. Although $n_{\rm Ser,bulge}$ changes more significantly, it plays a relatively less important role in determining CVC shape (see Section 3.3). As a result, these high $\sigma_{\rm e}$ galaxies exhibit nearly the same CVC shape.
4 Conclusions
In the seventh paper of the MaNGA DynPop project, we derive the CVCs (or intrinsic RCs) for 6000 nearby galaxies from stellar dynamical models in (Zhu et al., 2023b). The amplitude and shape of CVCs (corrected for the inclination angle and seeing effects) are closely related to the inner gravitational potential of galaxies, providing key insights into galaxy formation and evolution scenarios. By combining these with spatially resolved stellar population properties (Lu et al., 2023), we study the scaling relations between CVCs (shape and amplitude) and other galaxy properties. We propose a simple BDH model that includes a bulge, a disk and a dark matter halo to predict the shape of CVCs across different galaxy types.
The main conclusions are summarized as below.
  • The amplitude of CVCs, characterized by the circular velocity at the half-light radius, $V_{\rm circ}(R_{\rm e}^{\rm maj})$, or the maximum circular velocity within the FoV, $V_{\rm circ}^{\rm max}$, is linearly related to the velocity dispersion within one effective radius $\sigma_{\rm e}$, linking the Tully-Fisher relation (Tully & Fisher, 1977) and the Faber-Jackson relation (Faber & Jackson, 1976). The best-fitting relations are $V_{\rm circ}(R_{\rm e}^{\rm maj})\approx 1.65\times\sigma_{\rm e}$ and $V_{\rm circ}^{\rm max}\approx 1.75\times\sigma_{\rm e}$ with a small error of 9% or 8% (Figure 2), respectively.
  • The shape of CVCs (rising, flat, and declining) correlates with galaxy dynamical and stellar population properties: galaxies with declining CVCs are massive, early-type, early-quenched, old, metal-rich, and exhibit high $B/T$, low spin, high velocity dispersion, and flat age gradient, while galaxies with rising CVCs exhibit the opposite properties (Figure 3).
  • We propose a BDH model, which includes a Sersic bulge, an exponential disk, and a NFW dark matter halo, to quantify the shape of CVCs. In this model, given the bulge-to-total ratio $B/T$, the dark matter fraction within $f_{\rm DM}(<R_{\rm e})$, the Sersic index of the bulge component $n_{\rm Ser,bulge}$, the ratio between effective radii of bulge and disk $R_{\rm e,bulge}/R_{\rm e,disk}$, and the ratio between the scale radius of NFW halo and the luminous half-light radius $r_{\rm s}/R_{\rm e}$, one can predict the shape of CVCs. We test this model for galaxies within different $\sigma_{\rm e}$ bins, in which the free parameters are predicted from their empirical scaling relations with $\sigma_{\rm e}$ (see Figure 4 and figure 2 in Lu et al., 2024), finding a nearly unbiased consistency with JAM-derived CVC shapes (Figure 5). The BDH model quantitatively confirms that the shape of CVCs is mainly driven by $B/T$, $f_{\rm DM}(R_{\rm e)}$, and $n_{\rm Ser,bulge}$: galaxies with larger $B/T$, lower $f_{\rm DM}(R_{\rm e})$, and larger $n_{\rm Ser,bulge}$ tend to have declining CVCs, while rising CVCs are found in galaxies with opposite properties.
  • The evolution of CVC shapes on the mass-size plane is closely related to the evolution of $B/T$, $f_{\rm DM}(R_{\rm e})$, and $n_{\rm Ser,bulge}$, supporting the scenario of two evolutionary channels (e.g. Cappellari et al., 2013b; van Dokkum et al., 2015; Zhu et al., 2024): (i) in situ star formation (through gas accretion or gas-rich minor mergers) moving galaxies from left to right (CVC shape from rising to declining), inducing bulge growth, reducing central dark matter fraction, and increasing the Sersic index of bulge; (ii) dry mergers moving gas-poor galaxies along the constant $\sigma_{\rm e}$ lines upwards, mildly increasing $B/T$ and the Sersic index of bulge, leaving the central dark matter fraction unchanged, and finally leading to the same CVC shapes along the direction of the ZOE (Figure 6).
We will release all data derived from the stellar dynamical models, including 3D mass distributions (mass density profiles), 2D mass distributions (surface mass density maps), and circular velocity curves. One can use the MGE coefficients and best-fitting free parameters taken from the catalog (Zhu et al., 2023b) to derive the mass distributions of stellar, dark matter, and total components. The database and a Python script for the calculation will be provided on the website of MaNGA DynPop (https:manga-dynpop.github.io).
This work is supported by the National Science Foundation of China (Grant No. 11821303 to SM). KZ acknowledges the support from the Shuimu Tsinghua Scholar Program of Tsinghua University. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
Appendix A Dependence of the BDH model on the assumptions
We adopt a spherical exponential disk assumption, $V_{\rm disk}(r)=\sqrt{GM_{\rm disk}(r)/r}$, in the standard BDH model (Figure 5), which might not be a good approximation for LTGs. Here we test if the adopted assumptions affect the results. In Figure 7, we adopt the infinitely thin disk (Freeman, 1970) assumption for all galaxies (including ETGs and LTGs). Under this alternative extreme assumption, the circular velocity at radius $r$ (in the plane of disk) contributed from the disk component is
$V_{\rm disk}(r)=\frac{GM_{\rm\infty,disk}}{R_{\rm d}}2y^{2}\left[I_{0}(y)K_{0}(y)-I_{1}(y)K_{1}(y)\right],$ (A1)
where $M_{\rm\infty,disk}$ is the total mass of disk, $R_{\rm d}=R_{\rm e,disk}/1.678$ is the disk scale length, $y=(r/R_{\rm d})$, and $I_{\rm n}$ and $K_{\rm n}$ are the modified Bessel functions (Freeman, 1970). We find the predicted normalized $V_{\rm circ}$ profiles under the two extreme assumptions (spherical exponential disk or thin exponential disk) are nearly identical. In Figure 8, we fix the Sersic index of the bulge component to be 2.83, which is similar to Hernquist (1990) profile (see Appendix A and figure A1 in Vitral & Mamon, 2020), and find that adopting a universal Hernquist bulge can roughly predict the trend of CVCs (rising, flat or declining) but can not quantitatively reproduce the shape of CVCs in the inner region.
Figure 7: Similar to Figure 5, but adopt an alternative extreme assumption that all galaxies have infinitely thin exponential disk (Freeman, 1970).
Figure 8: Similar to Figure 5, but use a fixed $n_{\rm Ser,bulge}=2.83$ that corresponds to a Hernquist profile in the inner region (Appendix A and figure A1 in Vitral & Mamon, 2020). The trends of CVC shape changing with $\sigma_{\rm e}$ still exist, but BDH model predicted curves tend to overestimate at the low $\sigma_{\rm e}$ end and underestimate at the high $\sigma_{\rm e}$ end.